OpenStudy (anonymous):

HELP PLEASE The first four terms of a number sequence is given by: 3,5,7,9 Write down a. The 5th term b. The 10th term c.the nth term,in terms of n. (I got the first two (a,b) but i'm unable to get c)

OpenStudy (rock_mit182):

A sequence \[a_{1}, a _{2}, a_{_{3}} .. . .. .a _{n} \] is an arithmetic sequence if there is a real number d such that for every positive integer k, \[a _{k+1}=a _{k} + d \]

OpenStudy (anonymous):

common difference d=5-3=2 tn=a+(n-1)d here a=3 find tn

OpenStudy (anonymous):

Didnt understand

OpenStudy (anonymous):

Tn is 2n+1

OpenStudy (rock_mit182):

The number \[d = a _{k+1} - a _{k}\] is called the common difference of the sequence.

OpenStudy (anonymous):

tn=3+(n-1)2 =3+2n-2 =2n+1

OpenStudy (anonymous):

I just said that @surjithayer

OpenStudy (anonymous):

Ik and the common difference is 2 here. @rock_mit182

OpenStudy (rock_mit182):

of course and the first term is 3

OpenStudy (rock_mit182):

so you could use this formula

OpenStudy (anonymous):

Yeah so how am i going to find the nth term in the terms of n??

OpenStudy (anonymous):

i did and found out for the first two

OpenStudy (rock_mit182):

\[a _{n} = a _{1} + (n-1)d\]

OpenStudy (rock_mit182):

where \[a _{n}\] represents the nth term, that you want to find

OpenStudy (rock_mit182):

and n, also represents that nth term

OpenStudy (anonymous):

Okay i get it thanks:)

OpenStudy (rock_mit182):

;)